OpenStudy (anonymous):

the perimeter of a rectangle is 82m the length is 2m more than twice the width . what would the demenisions be then?

OpenStudy (anonymous):

let the width be x and length be 2x+2 perimeter=2(l+w) 82=2(2x+2+x) 3x+2=82/2 solve for x

OpenStudy (anonymous):

The width can be represented as W and the length can be represented as L = 2W + 2 The perimeter of the rectangle is 82m, and the formula for perimeter is P = 2L + 2W, So we can plug the numbers in and do the following ... P = 2L + 2W 82 = 2(2W + 2) + 2W Multiply 2 to what is inside the parenthesis 82 = 4W + 4 + 2W Subtract 4 on both sides and combine like terms (4W + 2W = 6W) 78 = 4W Divide by 6 on both sides W = 13 So we come to the conclusion that the length L is equal to 28 meters and the width W is equal to 13 meters. We can check out answer by using the perimeter formula. 2L + 2W 2(28) + 2(13) = 82 Our answer is complete.